OpenStudy (anonymous):

akar positi dari x^3-7x+6=0

OpenStudy (turingtest):

does that mean factor and solve?

OpenStudy (akshay_budhkar):

yes in malay

OpenStudy (turingtest):

oh it's cubic.. interesting, malay, I like the diversity here

OpenStudy (akshay_budhkar):

yea.. rather i speak 8 languages myself, so u know a very diverse person :P

myininaya (myininaya):

1-7+6=0

myininaya (myininaya):

one solution is therefore x=1

myininaya (myininaya):

use synthetic division to find the other two

myininaya (myininaya):

1 | 1 0 -7 6 | 1 1 -6 ----------------------- 1 1 -6 | 0 so solve \[x^2+x-6=0\] to find your other zeros

OpenStudy (anonymous):

positive roots?

OpenStudy (anonymous):

positeve root of x^3-7x+6=0?

myininaya (myininaya):

yes we have one positive root above x=1 now find the other by factoring x^2+x-6 and setting both of its factors =0

myininaya (myininaya):

hint: what is 3(-2) and 3+(-2)

OpenStudy (anonymous):

this is the opition a.1and2 b.1,2,and 3 c.1,2 and -3 d.-1 and 2 e.-1,2 and -3

OpenStudy (pokemon23):

turning test http://openstudy.com/study#/updates/4f0a576be4b014c09e647b5d